The adult maintenance dosage for phenobarbital as an antiepileptic can be as great as 3 mg/day for each kilogram of body mass. If the dose is administered by injecting a solution with a concentration of 130 mg/ml, how many milliliters are injected during the course of a day for an 86.67-kg patient?
Answer #1

ANSWER.

* Calculation of volume(ml) of drug required.

      Formula

         Volume(ml) = Desired dose/Dose in hand * Quantity.

    Here

          Ordered dose = 3mg/kg/day.

          Weight of the patient = 86.67 kg.

                Desired dose(D) = 3 * 86.67

                                            = 260.01 mg

                Dose in hand (H) = 130 mg.

                     Quantity(Q) = 1 ml.

       As per above formula

            Volume(ml) of drug required = D/H * Q

                                                          = 260.01 / 130 * 1

                                                          = 2 * 1

                                                          = 2 ml / day.